Find the equation of the line. Write it using function notation.
Through (8,-5) and perpendicular to 7y=x-14

Answer by ewatrrr(24785)

7y=x-14 0r y = (1/7)x - 2
...
New Line (Perpendicular) m = -7/1
y + 5 = -7(x-8)
y = -7x + 51
0r
7x + y = 51 in Standard Form